OpenStudy (anonymous):

What is the equation of a line that passes through (-6, 2) and has a slope of −12 ?

OpenStudy (godlovesme):

Use the point-slope formula which is y-y1=m(x-x1) Y1= 2 X1= -6 And m= -12 can u try? :) @tyetianna

OpenStudy (anonymous):

no